# Factorization in Finitely-Presented Monoids

### v1

@@ 5 Normalizing Monoids
It is easy to see that $\mathsf{N}(M)$ is a submonoid of $M$, and $M^{\times}\subseteq\mathsf{N}(M)$, for any monoid $M$. For this reason, $\mathsf{N}(M)$ is referred to as the normalizing submonoid of $M$. Given a finitely-generated monoid $M$, the submonoid $\mathsf{N}(M)$ need not be finitely-generated, though that is the case in certain situations–see, e.g., [32, Corollary 4.4.12]. Additionally, each normalizing monoid satisfies the left and right Ore conditions (i.e., $aM\cap bM\neq\emptyset$, and $Ma\cap Mb\neq\emptyset$, for all $a,b\in M$), and therefore each cancellative normalizing monoid has a quotient group–see [12, Theorems 1.24 and 1.25] for more details.
There is extensive literature on normal elements in the ideals and Factorization Theory of monoids and rings–see, e.g., [2, 34, 41, 14]. Normalizing Krull monoids, in particular, have received attention within the factorization context [4, 22].
Let us record another interesting fact about normal elements, which has been previously noted in more specialized contexts.
@@
Our next result is an analogue of the well-known fact that every atomic finitely-generated commutative cancellative monoid has accepted elasticity (in terms of atoms)–see, e.g. [18, Remark 3.11(5)]. Note that unlike the atomic commutative setting, we must impose the assumption that our monoid is BF, to exclude, for example, abelian groups–see the discussion in Section 3 for more details.
@@
By Dickson’s Theorem [24, Theorem 1.5.3], there are only finitely many minimal points in the image $f(S)$ of $f$, relative to the usual product order on $\mathbb{N}^{2n}$ (i.e., $(p_{1},p_{2},\dots,p_{2n})\preceq(r_{1},r_{2},\dots,r_{2n})$ if and only if $p_{i}\leq r_{i}$ for each $i$, in the standard ordering $\leq$ on $\mathbb{N}$). Let $T\subseteq S$ be the inverse image under $f$ of the set of minimal points in $f(S)$. Since $f$ is clearly injective, $T$ is finite, and so we can choose $(\nu(a),\nu(b))\in T$ such that $|\nu(a)|/|\nu(b)|$ is maximal. (Note that, by Lemma 4, if $c\in\langle X\rangle$ is such that $c=_{M}1$, then $c=1$. So $|\nu(a)|\neq 0\neq|\nu(b)|$, for all $(\nu(a),\nu(b))\in S$.) We claim that
@@
(2) According to [14, Theorem 5.8], every finitely-generated normalizing monoid satisfies the ascending chain condition on principal ideals. Moreover, according to [43, Corollary 2.6], an acyclic monoid $M$ is atomic if and only if $M$ has a generating set $X$ such that every ascending chain of principal ideals, starting with $MxM$ for some $x\in X$, stabilizes. Therefore the desired conclusion follows from (1).
@@
It can be shown that every monoid which is atomic, finitely-generated, commutative, and unit-cancellative has finite elasticity [15, Corollary 3.4] (in terms of atoms), but it need not be accepted [18, Remark 3.11(5)]. The example constructed in [18, Remark 3.11(5)] is reduced, and so, in view of Proposition 15(3), we cannot replace “cancellative” with “unit-cancellative” in the hypotheses of Proposition 13. We discuss in Section 6 the role of the normalizing assumption in Proposition 13 and Theorem 14.
@@ 6 Examples
Given a monoid $\langle X\mid R\rangle$, denote by $G_{l}(X\mid R)$, respectively $G_{r}(X\mid R)$, the undirected graph having vertex set $X$, and an edge between $x$ and $y$ (for $x,y\in X$) if and only if there exist $a,b\in\langle X\rangle$ with the property that $(xa,yb)\in R$, respectively $(ax,by)\in R$. A monoid $M$ is called Adyan if $M=\langle X\mid R\rangle$, where $X$ is finite, $(a,1),(1,a)\notin R$ for all $a\in\langle X\rangle$, and $G_{l}(X\mid R)$ and $G_{r}(X\mid R)$ are acyclic.
@@
Every Adyan semigroup embeds in a group, and is therefore cancellative.
In particular, a monoid of the form $\langle X\mid(a,b)\rangle$ is cancellative, provided that $a,b\in\langle X\rangle\setminus\{1\}$, and $a$ and $b$ begin, as well as end, with different elements of $X$. See [4, Section 2] for a more detailed discussion of the above result, [39] for a more general criterion, and [17] for another cancellativity criterion for monoids with presentations.
@@
If either $u$ does not end in $x$, or $v$ does not begin with $y$, then any applications of the defining relation to $a$, that results in a longer word, can also be applied to $(uy)z(xv)$, which contradicts the choice of $a$. On the other hand, if $u=u^{\prime}x$ and $v=yv^{\prime}$ for some $u^{\prime},v^{\prime}\in\langle x,y,z\rangle$, then
@@
and both admit the same expansions thereafter. Therefore we may assume that $a$ does not contain $yx$ as a subword.
@@
Now, let $a\in\langle u,v,x,y\rangle$ be such that $|a|=k$. If $a$ contains more than one instance of $u$ and $v$, then, in view of the defining relations, $\max(\mathsf{L}_{M}(a))\leq\rho_{k-2}(M)+3$. Thus if $m\in\mathcal{U}_{k}(M)$ is such that $m\geq\rho_{k-2}(M)+4$, then the difference between $m$ and the successive element of $\mathcal{U}_{k}(M)$ (if there is one) must be $2$, since it must be the result of applying repeatedly the relation $(xy,y^{n}x)$ to a shorter word. It follows that $\mu_{k}(M)\leq\rho_{k-2}(M)+4$. ∎
@@ 7 Summary
It is well-known that commutative finitely-generated cancellative monoids have accepted elasticity [24, Theorem 3.1.4], satisfy the Structure Theorem for Unions [18, Theorem 3.6], and are often fully elastic (e.g., [45, Theorem 1.2]), in terms of factorization into atoms. In this paper we prove analogous statements for certain non-commutative monoids, in the broader context of factorization into generators (Proposition 13, and Theorems 14 and 9), and give examples showing that, even for finitely-presented cancellative monoids, these properties need not always hold (Propositions 18 and 21). Much remains to be explored, however, as summarized in the following general problem.
Problem 22.
@@ Acknowledgements
This work was supported by the Austrian Science Fund FWF, Project P36852-N. Furthermore, the authors would like to thank Salvatore Tringali for fruitful discussions.

### v2

@@ 5 Normalizing Monoids
It is easy to see that $\mathsf{N}(M)$ is a submonoid of $M$, and $M^{\times}\subseteq\mathsf{N}(M)$, for any monoid $M$. For this reason, $\mathsf{N}(M)$ is referred to as the normalizing submonoid of $M$. Given a finitely-generated monoid $M$, the submonoid $\mathsf{N}(M)$ need not be finitely-generated, though that is the case in certain situations–see, e.g., [41, Corollary 4.4.12]. Additionally, each normalizing monoid satisfies the left and right Ore conditions (i.e., $aM\cap bM\neq\emptyset$, and $Ma\cap Mb\neq\emptyset$, for all $a,b\in M$), and therefore each cancellative normalizing monoid has a quotient group–see [16, Theorems 1.24 and 1.25] for more details.
There is extensive literature on normal elements in the Ideal Theory and Factorization Theory of monoids and rings–see, e.g., [2, 44, 52, 18]. Normalizing Krull monoids, in particular, have received attention within the factorization context [30, 4].
Let us record another interesting fact about normal elements, which has been previously noted in various other settings (see [30, 29]).
@@
Our next result is an analogue of the well-known fact that every atomic finitely-generated commutative cancellative monoid has accepted elasticity (in terms of atoms)–see, e.g. [24, Remark 3.11(5)]. Note that unlike the atomic commutative setting, we must impose the assumption that our monoid is BF, to exclude, for example, abelian groups–see the discussion in Section 3 for more details.
@@
By Dickson’s Theorem [32, Theorem 1.5.3], there are only finitely many minimal points in the image $f(S)$ of $f$, relative to the usual product order on $\mathbb{N}^{2n}$ (i.e., $(p_{1},p_{2},\dots,p_{2n})\preceq(r_{1},r_{2},\dots,r_{2n})$ if and only if $p_{i}\leq r_{i}$ for each $i$, in the standard ordering $\leq$ on $\mathbb{N}$). Let $T\subseteq S$ be the inverse image under $f$ of the set of minimal points in $f(S)$. Since $f$ is clearly injective, $T$ is finite, and so we can choose $(\nu(a),\nu(b))\in T$ such that $|\nu(a)|/|\nu(b)|$ is maximal. (Note that, by Lemma 5, if $c\in\langle X\rangle$ is such that $c=_{M}1$, then $c=1$. So $|\nu(a)|\neq 0\neq|\nu(b)|$, for all $(\nu(a),\nu(b))\in S$.) We claim that
@@
(2) According to [18, Theorem 5.8], every finitely-generated normalizing monoid satisfies the ascending chain condition on principal ideals. Moreover, according to [56, Corollary 2.6], an acyclic monoid $M$ is atomic if and only if $M$ has a generating set $X$ such that every ascending chain of principal ideals, starting with $MxM$ for some $x\in X$, stabilizes. Therefore the desired conclusion follows from (1).
@@
It can be shown that every monoid which is atomic, finitely-generated, commutative, and unit-cancellative has finite elasticity [19, Corollary 3.4] (in terms of atoms), but it need not be accepted [24, Remark 3.11(5)]. The example constructed in [24, Remark 3.11(5)] is reduced, and so, in view of Proposition 16(3), we cannot replace “cancellative” with “unit-cancellative” in the hypotheses of Proposition 14. We discuss in Section 6 the role of the normalizing assumption in Proposition 14 and Theorem 15.
@@ 6 Examples
Given a monoid $\langle X\mid R\rangle$, where $(a,1),(1,a)\notin R$ for all $a\in\langle X\rangle$, denote by $G_{l}(X\mid R)$, respectively $G_{r}(X\mid R)$, the undirected graph having vertex set $X$ and edge set $R$, such that for each $(a,b)\in R$ there is an edge between the initial letters of $a$ and $b$, respectively between the final letters of $a$ and $b$. A monoid $M$ is called Adyan if $M=\langle X\mid R\rangle$, where $X$ and $R$ are finite, $(a,1),(1,a)\notin R$ for all $a\in\langle X\rangle$, and $G_{l}(X\mid R)$ and $G_{r}(X\mid R)$ are acyclic.
@@
Every Adyan monoid embeds in a group, and is therefore cancellative.
In particular, a monoid of the form $\langle X\mid(a,b)\rangle$ is cancellative, provided that $a,b\in\langle X\rangle\setminus\{1\}$, and $a$ and $b$ begin, as well as end, with different elements of $X$. We note that our description of Adyan’s result is based on that given in [49], which also contains a more general criterion, though in that publication it is stated for semigroups rather than monoids. See [23] for another cancellativity criterion for monoids with presentations.
@@
If either $u$ cannot be expanded to a word that ends in $x$, or $v$ cannot be expanded to a word that ends in $y$, then any applications of the defining relation to $a$, that results in a longer word, can also be applied to $(uy)z(xv)$, which contradicts the choice of $a$. On the other hand, if $u=_{M}u^{\prime}x$ and $v=_{M}yv^{\prime}$ for some $u^{\prime},v^{\prime}\in\langle x,y,z\rangle$, then
@@
and the two admit the same expansions thereafter. Therefore we may assume that $a$ does not contain $yx$ as a subword.
@@
Now, let $a\in\langle u,v,x,y\rangle$ be such that $|a|=k$. If $a$ contains more than one instance of $u$ and $v$, then, in view of the defining relations, $\max(\mathsf{L}_{M}(a))\leq\rho_{k-2}(M)+3$. Thus if $m\in\mathcal{U}_{k}(M)$ is such that $m\geq\rho_{k-2}(M)+4$, then the difference between $m$ and the successive element of $\mathcal{U}_{k}(M)$ (if there is one) must be at least $2$, since it must be the result of applying repeatedly the relation $(xy,y^{n}x)$ to a shorter word. It follows that $\mu_{k}(M)\leq\rho_{k-2}(M)+4$. ∎
@@ 7 Summary
It is well-known that commutative finitely-generated cancellative monoids have accepted elasticity [32, Theorem 3.1.4], satisfy the Structure Theorem for Unions [24, Theorem 3.6], and are often fully elastic (e.g., [57, Theorem 1.2]), in terms of factorization into atoms. In this paper we prove analogous statements for certain non-commutative monoids, in the broader context of factorization into generators (Proposition 14, and Theorems 15 and 10), and give examples showing that, even for finitely-presented cancellative monoids, these properties need not always hold (Propositions 19 and 22). Much remains to be explored, however, as summarized in the following general problem.
Problem 23.
@@ Acknowledgements
This work was supported by the Austrian Science Fund FWF, Project P36852-N. Furthermore, the authors would like to thank Salvatore Tringali for fruitful discussions, and the anonymous referee for pointing us to relevant literature and providing helpful suggestions.